# What Do You Need for Diverse Trajectory Composition in Diffusion Planning?

### v2

@@ A.6.1 Broader Impacts
A.6.2 How Should We Select a Method for Composition?
@@
Hierarchical diffusion appears to be a strong way to incorporate both the benefits of long-horizon planning and short-horizon consistency. Besides the empirical benefits shown in the hierarchical diffusion planning papers discussed in 2, more evidence exists that hierarchical models can learn to minimize dynamics error in fewer training steps [19]. One downside of these methods is that how the task is hierarchically decomposed either needs to be set manually, requiring more tuning, or must rely on a more complex algorithm [42] which increases implementation cost and other tuning. Image models have generally shown that forgoing manually set hierarchy and instead letting models jointly learn both high level structure and low level detail can work well given proper data and architecture decisions.
@@
A.6.3 Explaining Composition in Other Diffusion Planning Works
@@
Hierarchical Diffusion Planning: As discussed in Section 2, several works approach the composition problem by separately generating sub-sequences with a diffusion model, and then composing them. Exactly how these low-level subsequences are "orchestrated" into a single coherent plan varies, but each exhibits aspects of locality and shift equivariance.
[10] propose Hierarchical Diffuser, which uses two separate diffusion models: a high level one which generates a "plan sketch", and a low level one which generates a sequence connecting the states in the plan sketch to "fill-in" the high-level plan. Functionally similar was the prior HDI paper from [42], which has a different form of generating sub-goals unrelated to diffusion. These approaches enable a form of shift equivariance by data augmentation: because the low-level subsequences generated by the low-level diffusion model are randomly cut out of the longer trajectories in the training data, they have their relative position within longer sequences removed. This also enables locality by disentangling the joint prediction of states within a particular subsequence to any other states outside of it, so each subsequence functions as a small local window without connectivity to previous or future states, preventing confounding from preventing stitching. [10] also confirms some of our findings that Diffuser with a larger kernel size often produces more coherent long-term plans (4.3) but that this trades-off with worse ODD compositional generalization (Appendix E).
[48] proposes a hierarchical method called CompDiffuser which, like the Hierarchical Diffuser paper, generates sub-sequences. Instead of having a separate high-level planner, they have each sub-sequence condition its generation on its immediate surrounding neighbour subsequence (3.2). This preserves shift equivariance in the same manner as Hierarchical Diffuser, but maintains locality through this process of having each sub-sequence only attend to its neighbours at each step, which stops causal confusion.
@@
Results from [9] shows that this approach can enable compositionality, but only when keeping short or no memory and only generating short plans, essentially using the diffusion forcing model to approximate short-horizon Model Predictive Control (E.2). When extending the generation length or memory, they find that it instead reproduces existing trajectories.
Follow up work from [62] extends Diffusion Forcing by replacing their RNN-based backbone with a transformer, and introducing the idea of combining the scores of the same model conditioned with several history lengths to try to achieve both long-sequence compositionality and short-term reactivity. They tested this approach on a task where a robot must take a fruit in a slot and move it to another slot. This task requires both memory (of which fruit the model should move) and short term reactivity (a human perturbs the robot mid-trajectory). However, the training data does not have both memory of the original fruit location and examples of recovery behaviour from the human perturbance, requiring composition of training sub-sequences. They found that the model that keeps the full trajectory history overfits to the training datasets showing the original fruit, while only using short-horizon models would compose recovery behaviour but forget the original task. Ultimately they found combining the score produced trajectories that both remembered important fast information and were capable of composing sub-trajectories for reactive control. This supports our argument highlighting the importance of locality for composition, but highlights that locality alone can run into issues in long-horizon tasks, as discussed in Section 6.1. For details, see Sections 6.4: Task 3, C.8, and D.4.
@@
This differs from the diffusion planning framework in that a long-horizon plan is not explicitly generated, so any long-horizon task guidance is obtained from conditioning the policy (like in goal-conditioned BC) rather than from explicit planning. Thus, in these models compositionality essentially comes from the same place that it does in TD learning: because no planning is done, the model composes through replanning. This replanning, like any explicit one-step or few-step policy, assumes the environment is Markovian and so does not include temporal information or preserve long history as conditioning, thus preserving shift equivariance and locality as we discussed in 3.2.

### v3

@@ VI-H1 Broader Impacts
VI-H2 How Should We Select a Method for Composition?
@@
Hierarchical diffusion appears to be a strong way to incorporate both the benefits of long-horizon planning and short-horizon consistency. Besides the empirical benefits shown in the hierarchical diffusion planning papers discussed in II, more evidence exists that hierarchical models can learn to minimize dynamics error in fewer training steps [20]. One downside of these methods is that how the task is hierarchically decomposed either needs to be set manually, requiring more tuning, or must rely on a more complex algorithm [50] which increases implementation cost and other tuning. Image models have generally shown that forgoing manually set hierarchy and instead letting models jointly learn both high level structure and low level detail can work well given proper data and architecture decisions.
@@
VI-H3 Explaining Composition in Other Diffusion Planning Works
@@
Hierarchical Diffusion Planning: As discussed in Section II, several works approach the composition problem by separately generating sub-sequences with a diffusion model, and then composing them. Exactly how these low-level subsequences are "orchestrated" into a single coherent plan varies, but each exhibits aspects of locality and shift equivariance.
Chen et al. [10] propose Hierarchical Diffuser, which uses two separate diffusion models: a high level one which generates a "plan sketch", and a low level one which generates a sequence connecting the states in the plan sketch to "fill-in" the high-level plan. Functionally similar was the prior HDI paper from Li et al. [50], which has a different form of generating sub-goals unrelated to diffusion. These approaches enable a form of shift equivariance by data augmentation: because the low-level subsequences generated by the low-level diffusion model are randomly cut out of the longer trajectories in the training data, they have their relative position within longer sequences removed. This also enables locality by disentangling the joint prediction of states within a particular subsequence to any other states outside of it, so each subsequence functions as a small local window without connectivity to previous or future states, preventing confounding from preventing stitching. Chen et al. [10] also confirms some of our findings that Diffuser with a larger kernel size often produces more coherent long-term plans (4.3) but that this trades-off with worse ODD compositional generalization (Appendix E).
Luo et al. [57] proposes a hierarchical method called CompDiffuser which, like the Hierarchical Diffuser paper, generates sub-sequences. Instead of having a separate high-level planner, they have each sub-sequence condition its generation on its immediate surrounding neighbour subsequence (3.2). This preserves shift equivariance in the same manner as Hierarchical Diffuser, but maintains locality through this process of having each sub-sequence only attend to its neighbours at each step, which stops causal confusion.
@@
Results from Chen et al. [9] shows that this approach can enable compositionality, but only when keeping short or no memory and only generating short plans, essentially using the diffusion forcing model to approximate short-horizon Model Predictive Control (E.2). When extending the generation length or memory, they find that it instead reproduces existing trajectories.
Follow up work from Song et al. [72] extends Diffusion Forcing by replacing their RNN-based backbone with a transformer, and introducing the idea of combining the scores of the same model conditioned with several history lengths to try to achieve both long-sequence compositionality and short-term reactivity. They tested this approach on a task where a robot must take a fruit in a slot and move it to another slot. This task requires both memory (of which fruit the model should move) and short term reactivity (a human perturbs the robot mid-trajectory). However, the training data does not have both memory of the original fruit location and examples of recovery behaviour from the human perturbance, requiring composition of training sub-sequences. They found that the model that keeps the full trajectory history overfits to the training datasets showing the original fruit, while only using short-horizon models would compose recovery behaviour but forget the original task. Ultimately they found combining the score produced trajectories that both remembered important fast information and were capable of composing sub-trajectories for reactive control. This supports our argument highlighting the importance of locality for composition, but highlights that locality alone can run into issues in long-horizon tasks, as discussed in Section V-A. For details, see Sections 6.4: Task 3, C.8, and D.4.
@@
This differs from the diffusion planning framework in that a long-horizon plan is not explicitly generated, so any long-horizon task guidance is obtained from conditioning the policy (like in goal-conditioned BC) rather than from explicit planning. Thus, in these models compositionality essentially comes from the same place that it does in TD learning: because no planning is done, the model composes through replanning. This replanning, like any explicit one-step or few-step policy, assumes the environment is Markovian and so does not include temporal information or preserve long history as conditioning, thus preserving shift equivariance and locality as we discussed in III-C.
@@
VI-I What explain Composition in Prior Monolithic Diffusion Planners?
Our results shown in Section III-A and throughout the paper indicate that diffusion planners, when used to generate long-horizon plans, tend to memorize exact sequences. This seems to contradict the results of Janner et al. [40] and Ajay et al. [1] which both also use monolithic planners. We suspect the following reasons are at play:
Most of their tasks are trained on the order of thousands of trajectories, which we find is usually sufficent to generalize.
The D4RL Maze datasets have been found not to require stitching [27], potentially letting the diffusion planner "cheat" on the task by simply memorizing appropriate trajectories.
For many of their experiments they use a shorter planning horizon of , allowing some stitching from replanning like with diffusion policies [13].